OpenStudy (anonymous):

4x+9-14x=3(4x-7) help solve please, with steps

OpenStudy (anonymous):

Alright, we have: 4x + 9 -14x = 3(4x-7) Let's start by distributing that 3. What would the equation look like after doing that?

OpenStudy (anonymous):

4x+9-14x=12x-21

OpenStudy (anonymous):

4x+9-14x=3(4x-7) first distribute the 3 4x+9-14x=12x-21 then add like terms -10x+9=12x-21 add 21 to both sides -10x+30=12x +10x +10x 30=22x x=30/22

OpenStudy (anonymous):

Sweet! Yep, okay, now combine like terms. What's it look like after doing that?

OpenStudy (anonymous):

thanks guys!